OpenStudy (anonymous):

what fraction is equivalent to 0.05 ??

OpenStudy (anonymous):

1/20

ganeshie8 (ganeshie8):

1/2 = 0.5 therefore 1/20 = 0.05

OpenStudy (anonymous):

to find it, if it is not obvious, start with \[\frac{5}{100}\]and then reduce

OpenStudy (anonymous):

lol there is more that one way to skin a cat

OpenStudy (anonymous):

Thx